If one inch on a map of Texas is 55 miles how many inches would the length of Texas (790 miles) be on the map. round to the nearest tenth of an inch
I got:
790*55=43450/12=3620.833=3620.9in

Answer by Theo(13342)

1 inch on the map is equal to 55 miles in the real world.

set up a ratio.

1 inch on the map divided by 55 miles in the real world = x inches on the map divided by y miles in the real world.

the ratio becomes 1/55 = x/y.

when y = 790, this ratio becomes 1/55 = x/790

solve for x to get x = 790 / 55 = 14.3636363636... inches on the map.